We have to solve the problem using Linear Equations. Here's the question... "All 28 members in Crestview High School's Ski Club went on a one-day ski trip. Members can rent skies for$16.00 per day or snow boards for $19.00 per day. The club paid a total of $478 for rental equipment. How many members rented skies & how many rented snowboards?
So far all I have is x= # of members rented skies ; y= # of members rented a snowboard
x+y=478
16x+19y=

Answer by macston(5194)

.
x=number of snowboards; y=number of skis
x+y=28
x=28-y Use this to substitute for x.
.
$19x+$16y=$478 Substitute for x from above.
$19(28-y)+$16y=$478
$532-$19y+$16y=$478
-$3y=-$54
y=18 ANSWER 1: 18 members rented skis.
x=28-y=28-18=10 ANSWER 2: 10 members rented snowboards.
.
CHECK:
$19x+$16y=$478
$19(10)+$16(18)=$478
$190+$288=$478
$478=$478